# Advancing Auto-Regressive Continuation for Video Frames

### v1

@@ Abstract
Recent advances in auto-regressive large language models (LLMs) have shown their potential in generating high-quality text, inspiring researchers to apply them to image and video generation. This paper explores the application of LLMs to video continuation, a task essential for building world models and predicting future frames. In this paper, we tackle challenges including preventing degeneration in long-term frame generation and enhancing the quality of generated images. We design a scheme named ARCON, which involves training our model to alternately generate semantic tokens and RGB tokens, enabling the LLM to explicitly learn and predict the high-level structural information of the video. We find high consistency in the RGB images and semantic maps generated without special design. Moreover, we employ an optical flow-based texture stitching method to enhance the visual quality of the generated videos. Quantitative and qualitative experiments in autonomous driving scenarios demonstrate our model can consistently generate long videos.
@@ 1 Introduction
Many researchers believe that LLMs have advantages in certain aspects, such as diverse generation, multi-modal fusion, and scalable model capability. Researchers aim to further harness the capabilities of the next-token prediction paradigm and explore scaling up to even larger models. LVM [3] demonstrates that auto-regressive learning solely on tokenized image and video frame sequences can yield reasonable scaling performance and some non-trivial next-frame inference results. WorldGPT [19] acquires an understanding of world dynamics by analyzing millions of videos across various domains. Some cutting-edge works, such as VideoPoet [30] and GAIA-1 [25], have already achieved impressive results in token-based video generation.
These works inspire us to explore the potential of LLMs further in video continuation. The significant computational power and engineering investment required make exploration and ablation experiments challenging for LLM-based visual generation. There are currently many open problems that have not been fully researched:
@@
    We establish a video continuation model based on a visual tokenizer and LLM architecture that has the potential to achieve emergent capabilities.
@@ 2 Related Work
Supervised data is often expensive and limited in quantity, leading many researchers to focus on visual pre-training [24, 4]. This approach harnesses large models and unlabeled data to mirror the emergent abilities of LLMs, enabling a generalized model to tackle various visual tasks. iGPT [12] pioneers using next-pixel prediction to train an auto-regressive model for image generation, but struggled with high-resolution images due to direct pixel distribution learning. ViT [13] transforms visual tasks into sequence modeling by dividing images into patches and extracting embeddings, showcasing the transformer’s potential for computer vision, albeit limited to image classification. Subsequent works [5, 74, 24] adopt similar patch-based approaches, employing masked image modeling (MIM) for pre-training. They utilize visual tokenizers based on VQ-VAE [60], learned via auto-encoding, to convert image patches into tokens. During pre-training, masked patch tokens are predicted to reconstruct the original image. The Emu series [51, 52, 66] introduce a cross-modal tokenizer to uniformly encode text, images, and videos into tokens for large multi-modal model pre-training.
@@ 3.1 Background
We define the video continuation task as generating the future frames $\{\tilde{I}_{t+1},\tilde{I}_{t+2},\tilde{I}_{t+3},\ ...\}$ given a sequence of past $t$ frames $\{I_{i}\in\mathbb{R}^{h\times w\times 3}|_{i=1,...,t}\}$. The inputs of our video continuation model are the three consecutive frames $I_{t-2}$, $I_{t-1}$, and $I_{t}$. Our auto-regressive paradigm allows us to generate an arbitrary number of frames iteratively. Our ARCON approach can generate creative minute-level videos, based on the model design paradigm of generating video frames in an auto-regressive manner. Our model comprises three decoupling components: (1) an image tokenizer encoding images and semantic maps into discrete tokens, (2) an LLM trained with the next token prediction task to perform the video continuation task in an auto-regressive manner, and (3) an image decoder that can decode discrete tokens to images.
@@
LLM.
Based on MAGVIT-v2’s token factorization technique [83], instead of predicting using a codebook of size $2^{18}$, we can predict using two concatenated codebooks, each of size $2^{9}$. The final result is that each $112\times 112\times 3$ frame will be converted into $784$ 1D tokens with a vocabulary size of $2^{9}$. We adopt the LLM architecture of LLaMA [57] which is a popular open-source model. We train the model from scratch with a context length of $16,384$ tokens, which can fit no more than $20$ images under MAGVIT-v2 [83] tokenizer. We package the data into a question-and-answer format and use a system prompt to specify whether the task is to continue with pure RGB tokens or to interleave semantic/RGB tokens. Once frames can be represented as token sequences, we can train the model to minimize the cross-entropy loss for predicting the next token.
@@ 3.2 Interleaving generation
The benefit of the auto-regressive framework is that multimodal data from a variety of data structures (text, images, videos, audio, etc.) can all be converted into discrete tokens so that language models can understand the information from the various modalities based on different codebooks. Many works have attempted to utilize this approach for translation between multimodal data [3, 30], or for image generation using auxiliary modalities (depth maps, segmentation maps, sketches, etc.) as control signals [80]. Auxiliary modalities can carry much structural information of an image at an information density much lower than that of an RGB image. We believe that generating sequences of auxiliary modal tokens along with sequences of RGB tokens helps the LLM simplify the video continuation task. By incorporating the task of continuing the auxiliary modal sequences, we essentially break down the video continuation task into two sub-tasks: continuing the auxiliary modality and translating between modalities. This decomposition allows the model to capture structural details with fewer tokens, while also enabling the generation of new objects and scenes that were not present in the input sequence. This approach reduces the computational cost for the model and enhances both the temporal consistency and creative flexibility of the video continuation.
The structure of our ARCON model is shown in Fig. 2. First, we extract the semantic segmentation maps of the video data using a pre-trained Uniformer [31]. Then we utilize an image tokenizer that can encode images from RGB pixel space to discrete token space. We use this tokenizer to extract RGB tokens from video frames, as well as semantic tokens from the corresponding auxiliary modalities. Finally, we employ an LLM based on LLaMA [57], in its training and inference phases, we interleave the semantic tokens and the RGB tokens, so that the model generates semantic tokens and then RGB tokens at each timestamp to achieve the video continuation task.
@@ 3.3 Token decoder
Representing high-definition images with very few discrete tokens inevitably leads to quality loss [83, 50]. To make up for this loss, one approach is to hope that the decoder can remember various texture details. Previous work [25, 35] has demonstrated that training a large decoder, such as a video diffusion decoder, can enhance visual quality. In our work, we investigate the use of a simple method inspired by reference-based super resolution [88, 37] to borrow texture information from high-resolution input frames, thereby enhancing video quality. We train the flow-based feature warping method on BDD100K using open-MAGVIT-v2 scheme [39]. We keep the encoder frozen and fine-tune the MAGVIT-v2 decoder. This fine-tuning approach enhances the quality of reconstructed videos in an offline manner.
Fig. 3 is a simple schematic diagram illustrating our decoder. A simple flow model predicts a flow f that aligns between two feature maps using several layers of $3\times 3$ convolution layers. We use the reference tokens from high-resolution input frames with the shape of $224\times 224\times 3$ during inference.
@@ 4 Experiments
This section conducts experiments on our discrete token-based auto-regressive model and verifies the suitability of the proposed image tokenizer and LLM for long-term video continuation tasks. Quantitative and qualitative results demonstrate that semantic tokens can help generate more creative and temporally consistent long videos.
@@ 4.2 Probing Experiments
We find that interleaving generated semantic maps and RGB images have a high degree of consistency, results shown in Fig. 5. We use the same semantic segmentation model to re-extract the semantic segmentation maps for the generated images and find that the accuracy of the semantic segmentation maps generated by the auto-regressive model can reach 76% on average on nuScenes validation set.
@@
Table 2: Comparison with other methods on nuScenes validation set. We achieve the best FVD results on the nuScenes dataset without the need for fine-tuning [8].
@@ 4.3 Quantitative results
We evaluate our ARCON model’s capability on video continuation and compare ARCON with related video generation methods in autonomous driving scenarios. As exhibited in Tab. 2, the quantitative results show that even without training or fine-tuning on the nuScenes dataset [8], our model still outperforms other baselines on Fréchet Video Distance (FVD) scores [58], which proves that our model is not only capable of generating high-quality videos but also has strong generalization ability.
@@ 4.4 Qualitative results
Our model possesses the capability to continuously generate scenes from the perspective of a moving vehicle, effectively simulating vehicle turns and seamlessly generating new scenes. Notably, the model exhibits an autonomous learning process, acquiring fundamental traffic knowledge. For instance, it rarely tries to turn left in the right turn lane, and demonstrates a tendency to decelerate or maneuver around vehicles positioned in front of it. Furthermore, the model consistently generates coherent traffic lights, lane markings, signs, and crosswalks within urban contexts. The outcomes of these capabilities are visually represented in Fig. 6.
@@ 4.5 Ablation study
As shown in Tab. 3, we investigate the efficacy of our simple flow-based feature warping method in the decoding process. We report the FVD metric on the in-distribution BDD100K [81] and out-of-distribution K700 datasets. Cross-frame feature transfer relies on the similarity between frames, with lower similarity increasing the difficulty of the transfer. We report the FVD metric at various frame rates, which result in differing inter-frame similarities. FVD metrics are evaluated on $400$ samples in validation split with $16$-frame clips at the resolution of $224\times 224$. The method has demonstrated benefits in both in-distribution scenarios and out-of-distribution scenarios.
@@
Table 3: FVD metrics on BDD100K [81] dataset and K700 dataset at different frame-rates. Under higher frame rate tests, the magnitude of FVD will increase, which, although not intuitive, can be corroborated by the metrics of related works [59, 78].
@@ 5 Conclusion
We develop an ARCON scheme that interleaves the generation of RGB tokens and semantic tokens. We adhere to a generative paradigm that separates structure from texture, demonstrating the semantic effect of semantic tokens on the generation of RGB tokens. Our approach can be improved at several aspects. At present, the representation of a single image necessitates hundreds of tokens, combined with the extensive parameter count of LLM, leading to notably slow inference speeds for each generated image. Novel encoding strategies, as proposed in recent studies [86, 16], can offer enhancements to the efficiency of this approach. Additionally, our texture transfer method lacks effective guidance for long-term generation tasks. Future research could investigate the integration of flow-based techniques with generative methods at the decoding stage [40]. Furthermore, a deeper exploration into the physical consistency of generative models is crucial, particularly in scenarios akin to potential collision situations, which are pivotal for autonomous driving applications.

### v3

# ARCON: Advancing Auto-Regressive Continuation for Driving Videos

@@ Abstract
Recent advancements in auto-regressive large language models (LLMs) have led to their application in video generation. This paper explores the use of Large Vision Models (LVMs) for video continuation, a task essential for building world models and predicting future frames. We introduce ARCON, a scheme that alternates between generating semantic and RGB tokens, allowing the LVM to explicitly learn high-level structural video information. We find high consistency in the RGB images and semantic maps generated without special design. Moreover, we employ an optical flow-based texture stitching method to enhance visual quality. Experiments in autonomous driving scenarios show that our model can consistently generate long videos.
@@ 1 Introduction
Many researchers believe that LLMs have advantages in certain aspects, such as diverse generation, multi-modal fusion, and scalable model capability. Researchers aim to further harness the capabilities of the next-token prediction paradigm and explore scaling up to even larger models. Sequential Large Vision Model (LVM) [2] demonstrates that auto-regressive learning solely on tokenized image and video frame sequences can yield reasonable scaling performance and some non-trivial next-frame inference results. WorldGPT [18] acquires an understanding of world dynamics by analyzing millions of videos across various domains. Some cutting-edge works, such as VideoPoet [32] and GAIA-1 [25], have already achieved impressive results in token-based video generation.
These works inspire us to explore the potential of LVMs further in video continuation. The significant computational power and engineering investment required make exploration and ablation experiments challenging for LVM-based visual generation. There are currently many open problems that have not been fully researched:
@@
    We establish a video continuation model based on a visual tokenizer and LVM architecture that has the potential to achieve emergent capabilities.
@@ 2 Related Work
Supervised data is often expensive and limited in quantity, leading many researchers to focus on visual pre-training [23, 3]. This approach harnesses large models and unlabeled data to mirror the emergent abilities of LLMs, enabling a generalized model to tackle various visual tasks. ImageGPT [11] pioneers using next-pixel prediction to train an auto-regressive model for image generation, but struggled with high-resolution images due to direct pixel distribution learning. ViT [12] transforms visual tasks into sequence modeling by dividing images into patches and extracting embeddings, showcasing the transformer’s potential for computer vision, albeit limited to image classification. Subsequent works [4, 76, 23] adopt similar patch-based approaches, employing masked image modeling (MIM) for pre-training. They utilize visual tokenizers based on VQ-VAE [62], learned via auto-encoding, to convert image patches into tokens. During pre-training, masked patch tokens are predicted to reconstruct the original image. The Emu series [53, 54, 68] introduce a cross-modal tokenizer to uniformly encode text, images, and videos into tokens for large multi-modal model pre-training.
@@ 3.1 Background
We define the video continuation task as generating the future frames $\{\tilde{I}_{t+1},\tilde{I}_{t+2},\tilde{I}_{t+3},\ ...\}$ given a sequence of past $t$ frames $\{I_{i}\in\mathbb{R}^{h\times w\times 3}|_{i=1,...,t}\}$. The inputs of our video continuation model are the three consecutive frames $I_{t-2}$, $I_{t-1}$, and $I_{t}$. Our auto-regressive paradigm allows us to generate an arbitrary number of frames iteratively. Our ARCON approach can generate creative minute-level videos, based on the model design paradigm of generating video frames in an auto-regressive manner. Our model comprises three decoupling components: (1) an image tokenizer encoding images and semantic maps into discrete tokens, (2) an LVM trained with the next token prediction task to perform the video continuation task in an auto-regressive manner, and (3) an image decoder that can decode discrete tokens to images.
@@
LVM.
Based on MAGVIT-v2’s token factorization technique [85], instead of predicting using a codebook of size $2^{18}$, we can predict using two concatenated codebooks, each of size $2^{9}$. The final result is that each $112\times 112\times 3$ frame will be converted into $784$ 1D tokens with a vocabulary size of $2^{9}$. We adopt the LLaMA [59] architecture which is a popular open-source model. We train the model from scratch with a context length of $16,384$ tokens, which can fit no more than $20$ images under MAGVIT-v2 [85] tokenizer. We package the data into a question-and-answer format and use a system prompt to specify whether the task is to continue with pure RGB tokens or to interleave semantic/RGB tokens. Once frames can be represented as token sequences, we can train the model to minimize the cross-entropy loss for predicting the next token.
@@ 3.2 Interleaving generation
The benefit of the auto-regressive framework is that multimodal data from a variety of data structures (text, images, videos, audio, etc.) can all be converted into discrete tokens so that language models can understand the information from the various modalities based on different codebooks. Many works have attempted to utilize this approach for translation between multimodal data [2, 32], or for image generation using auxiliary modalities (depth maps, segmentation maps, sketches, etc.) as control signals [82]. Auxiliary modalities can carry much structural information of an image at an information density much lower than that of an RGB image. We believe that generating sequences of auxiliary modal tokens along with sequences of RGB tokens helps the LVM simplify the video continuation task. By incorporating the task of continuing the auxiliary modal sequences, we essentially break down the video continuation task into two sub-tasks: continuing the auxiliary modality and translating between modalities. This decomposition allows the model to capture structural details with fewer tokens, while also enabling the generation of new objects and scenes that were not present in the input sequence. This approach reduces the computational cost for the model and enhances both the temporal consistency and creative flexibility of the video continuation. Similar ideas have also demonstrated effectiveness in recent speech generation work, Step-Audio [28].
The structure of our ARCON model is shown in Fig. 2. First, we extract the semantic segmentation maps of the video data using a pre-trained Uniformer [33]. Then we utilize an image tokenizer that can encode images from RGB pixel space to discrete token space. We use this tokenizer to extract RGB tokens from video frames, as well as semantic tokens from the corresponding auxiliary modalities. Finally, we employ an LVM based on LLaMA [59], in its training and inference phases, we interleave the semantic tokens and the RGB tokens, so that the model generates semantic tokens and then RGB tokens at each timestamp to achieve the video continuation task.
@@ 3.3 Token decoder
Representing high-definition images with very few discrete tokens inevitably leads to quality loss [85, 52]. To mitigate this, previous work [25, 37] has demonstrated that training a large decoder, such as a video diffusion decoder, can remember various texture details. In our work, we investigate the use of a simple method inspired by reference-based super resolution [90, 39] to borrow texture information from high-resolution input frames, thereby enhancing video quality. We train the flow-based feature warping method on BDD100K using open-MAGVIT-v2 scheme [41]. We keep the encoder frozen and fine-tune the MAGVIT-v2 decoder. This fine-tuning approach enhances the quality of reconstructed videos in an offline manner.
Fig. 3 is a schematic diagram illustrating our decoder. A simple flow model predicts a flow f that aligns between two feature maps using several layers of $3\times 3$ convolution layers. We use the reference tokens from high-resolution input frames with the shape of $224\times 224\times 3$ during inference.
@@ 4 Experiments
This section conducts experiments on our discrete token-based auto-regressive model and verifies the suitability of the proposed image tokenizer and LVM for long-term video continuation tasks. Quantitative and qualitative results demonstrate that semantic tokens can help generate more creative and temporally consistent long videos.
@@ 4.2 Probing Experiments
We find that interleaving generated semantic maps and RGB images have a high degree of consistency, results shown in Fig. 5. We use the same semantic segmentation model to re-extract the semantic segmentation maps for the generated images and find that the accuracy of the semantic segmentation maps generated by the auto-regressive model can reach 77.4% on average on nuScenes validation set.
@@
Table 2: Comparison with other methods on nuScenes validation set. $\dagger$ denotes general video generation algorithms. We achieve the best FVD results on the nuScenes dataset without the need for fine-tuning [7].
@@ 4.3 Quantitative results
We evaluate our ARCON model’s capability on video continuation and compare with related video generation methods. As exhibited in Tab. 2, the quantitative results show that even without training or fine-tuning on the nuScenes dataset [7], our model still outperforms other baselines on Fréchet Video Distance (FVD) scores [60], which proves that our model is not only capable of generating high-quality videos but also has strong generalization ability.
@@ 4.4 Qualitative results
Our model possesses the capability to continuously generate scenes from the perspective of a moving vehicle. Notably, the model exhibits an autonomous learning process, acquiring fundamental traffic knowledge. For instance, it rarely tries to turn left in the right turn lane, and demonstrates a tendency to decelerate or maneuver around vehicles positioned in front of it. Furthermore, the model consistently generates coherent traffic lights, lane markings, signs, and crosswalks within urban contexts. The outcomes of these capabilities are visually represented in Fig. 6.
@@ 4.5 Ablation study
As shown in Tab. 3, we investigate the efficacy of our simple flow-based feature warping method in the decoding process. We report the FVD metric on the in-distribution BDD100K [83] and out-of-distribution K700 [9] datasets. Cross-frame feature transfer relies on the similarity between frames, with lower similarity increasing the difficulty of the transfer. We report the FVD metric at various frame rates, which result in differing inter-frame similarities. FVD metrics are evaluated on $400$ samples in validation split with $16$-frame clips at the resolution of $224\times 224$. The method has demonstrated benefits in both in-distribution scenarios and out-of-distribution scenarios.
@@
Table 3: FVD metrics on BDD100K [83] dataset and K700 [9] dataset at different frame-rates. Under higher frame rate tests, the magnitude of FVD will increase, which, although not intuitive, can be corroborated by the metrics of related works [61, 80].
@@ 5 Conclusion
We develop an ARCON scheme that interleaves the generation of RGB tokens and semantic tokens. We adhere to a generative paradigm that separates structure from texture, demonstrating the effect of semantic tokens on the generation of RGB tokens. Our approach can be improved at several aspects. At present, the representation of a single image necessitates hundreds of tokens, combined with the extensive parameter count of LVM, leading to notably slow inference speeds. Novel encoding strategies [88, 15] can offer enhancements to the efficiency of this approach. Additionally, our texture transfer method lacks effective guidance for long-term generation tasks. Future research could investigate the integration of flow-based techniques with generative methods at the decoding stage [42]. Furthermore, exploration into the physical consistency of generative models is pivotal for autonomous driving applications.